Carbaugh Corporation has provided the following production and average cost data for two levels of monthly production volume. The company produces a single product.

  Production volume 4,300 units 5,300 units
  Direct materials $78.90 per unit $78.90 per unit
  Direct labor $52.20 per unit $52.20 per unit
  Manufacturing overhead $77.30 per unit $68.30 per unit

The best estimate of the total cost to manufacture 5,000 units is closest to: (Do not round intermediate calculations.)

$1,026,490

$1,015,315

$1,008,610

$977,320
Answer #1

NOTE:Variable cost per unit and Total fixed costs do not change with change in units.

Total manufacturing overhead for 4300 units=(4300*77.3)=$332,390

Total manufacturing overhead for 5300 units=(5300*68.3)=$361990

Variable manufacturing overhead per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(361990-332390)/(5300-4300)=$29.6 per unit

Fixed manufacturing overhead=361990-(29.6*5300)=$205110

Hence total cost for 5000 units=Fixed cost+Variable cost

=205110+(78.9+52.2+29.6)*5000 units

=205110+803500

=$1,008,610